# Maintaining Plasticity in Reinforcement Learning: A Cost-Aware Framework for Aerial Robot Control in Non-stationary Environments

### v1

@@ Abstract
Reinforcement learning (RL) has demonstrated the ability to maintain the plasticity of the policy throughout short-term training in aerial robot control. However, these policies have been shown to loss of plasticity when extended to long-term learning in non-stationary environments. For example, the standard proximal policy optimization (PPO) policy is observed to collapse in long-term training settings and lead to significant control performance degradation. To address this problem, this work proposes a cost-aware framework that uses a retrospective cost mechanism (RECOM) to balance rewards and losses in RL training with a non-stationary environment. Using a cost gradient relation between rewards and losses, our framework dynamically updates the learning rate to actively train the control policy in a disturbed wind environment. Our experimental results show that our framework learned a policy for the hovering task without policy collapse in variable wind conditions and has a successful result of 11.29% less dormant units than L2 regularization with PPO. Project website: https://aerialroboticsgroup.github.io/rl-plasticity-project/
@@ II-B Cognitive Neuroscience Inspirations in RL
In the context of RL, a series of studies inspired by neuroscience and cognitive science have been conducted. In [11], inspired by cognitive science, a relation to a primacy bias is established, which is a behaviour of relying on early interactions and ignoring the useful feature encountered later. It has been shown that a solution to this problem can be provided by periodically resetting the last layer of the network. In [12], valuable insights into human cognition and behaviour are highlighted by making intriguing connections between fast and slow learning processes. At the same time, it is also expressed in the permeability between RL and studies of human cognition and behaviour. According to the connections between neuroscience and RL, [13] proposes that the brain optimizes multiple cost functions to facilitate targeted learning, using efficient architectures tailored to specific cognitive tasks. It is thought that by combining both neuroscience and deep learning principles, the gap between biological and artificial intelligence can be closed according to this perspective [14]. Therefore, we can take inspiration from the relationship between reward-prediction circuits and error-processing systems to migrate plasticity loss in RL.
@@ IV-A Simulation Training and Evaluation
We evaluated hovering learning strategies learned by the PPO agent with wind perturbations. To investigate the loss of plasticity and maintaining plasticity strategies, we trained three different RL agents, including standard PPO, L2 regularization with PPO and RECOM with L2 PPO. In Fig. 3, we have shown the average rewards for the three agents, averaged over 5 different experiments performed at 20 million timesteps.
@@
During the policy training for long periods, it was observed that the L2 and RECOM approaches produced successful results. When we looked at the change in dormant units with the L2 and RECOM approaches, we observed similar changes. However, the L2 regularization with PPO was shown to have 2.42% more dormant units than RECOM with L2 PPO. This observation led us to investigate the phenomena of dormant units in different conditions.
@@ IV-B Validation of Learned Policies
In Table III, we used the mean square error (MSE) metric to evaluate the position error of different RL agents. The standard PPO achieved a success rate of only 30% in a non-stationary environment setting, with relatively large tracking errors in three axes due to policy collapse. In contrast, L2 regularization with PPO substantially improved the success rate to 88% while also significantly reducing the overall position tracking error compared with standard PPO. Our proposed RECOM with L2 PPO method demonstrated the highest success rate at 90% and the smallest horizontal errors. There are also linear MPC approaches with sub-centimeter performance, as shown in [21]. However, the goal of this paper is not only to propose a new controller for aerial robotics applications, but also to show that a mechanism that balances rewards and losses, inspired by neuroscience and cognitive science, can evolve a policy for hovering tasks without policy collapse in long-term training. An example of successful hovering under wind disturbance was visualized in Fig. 7.
@@ IV-C Discussions
Our experiments have carefully evaluated the proposed RECOM with L2 PPO framework, focusing on the long-term training performance of the aerial robot for the hovering task. According to the result, we show that the RECOM with L2 PPO both prevents policy collapse compared to standard PPO and organizes dormant units better than L2 regularization with PPO. While the standard PPO experienced a dramatic policy collapse after 10 million timesteps, RECOM with L2 PPO continued to attempt to maximize reward in the long-term training of the hovering task. In addition to achieve these gains, it showed that under the wind disturbance a validation performance of 90% according to the results in the RotorPy simulation environment. To better understand the contribution of rewards and losses to RL and their impact on policy stability, we considered insights from cognitive science and neuroscience.
@@
In standard RL formulations, these two signals, reward maximization and loss minimization, are usually handled indirectly (e.g., through a single scalar reward plus a training loss objective). However, by explicitly balancing the role of reward and loss in one retrospective cost term, RECOM is designed to emulate a richer feedback loop by evoking how biological systems adapt behaviour. Specifically, RECOM’s retrospective cost $\mathbb{C}_{\mathrm{ret}}$ combines the negative of cumulative reward $-\mathbb{R}[t]$ with the training loss $\mathbb{L}[t]$ into a single scalar. Therefore, it models the interplay between dopaminergic “go” signals (reward) and cortical/subcortical “caution” or “error” signals (loss). This unified feedback drives an adaptive update to the PPO learning rate, thereby aiding to preserve the agent’s plasticity over long horizons and contributing to preventing policy collapse. By periodically re-evaluating the learning rate based on recent rewards and losses, the RECOM allows the agent to both avoid policy collapse and reactivate dormant units more successfully and to adapt to changing environmental conditions throughout long-term training. Additionally, the RECOM update acts much like a negative feedback control loop, lowering the learning rate if the recent cost rises and increasing it if the cost drops. This would be align with how neural circuits modulate synaptic plasticity in response to aggregated outcomes: reward signals can boost learning rates, while substantial errors can dampen them to maintain stability.
@@ V Conclusion
In this work, we proposed a retrospective cost mechanism (RECOM) to balance rewards and losses specifically focusing on the aerial robot for hovering tasks in RotorPy with non-stationary environment scenarios. With the proposed RECOM with L2 PPO, policy collapse is prevented in long-term training and has 11.29% less dormant units in the policy network than L2 regularization with PPO at 20 million timesteps. The experimental results demonstrated that rewards and losses could be balanced with a retrospective mechanism inspired by the perspective of neuroscience and cognitive science.

### v2

@@ Abstract
Reinforcement learning (RL) has demonstrated the ability to maintain the plasticity of the policy throughout short-term training in aerial robot control. However, these policies have been shown to loss of plasticity when extended to long-term learning in non-stationary environments. For example, the standard proximal policy optimization (PPO) policy is observed to collapse in long-term training settings and lead to significant control performance degradation. To address this problem, this work proposes a cost-aware framework that uses a retrospective cost mechanism (RECOM) to balance rewards and losses in RL training with a non-stationary environment. Using a cost gradient relation between rewards and losses, our framework dynamically updates the learning rate to actively train the control policy in a variable wind environment. Our experimental results show that our framework learned a policy for the hovering task without policy collapse in variable wind conditions and has a successful result of 11.29% less dormant units than L2 regularization with PPO. Project website: https://aerialroboticsgroup.github.io/rl-plasticity-project/
@@ II-B Cognitive Neuroscience Inspirations in RL
In the context of RL, a series of studies inspired by neuroscience and cognitive science have been conducted. In [11], inspired by cognitive science, a relation to a primacy bias is established, which is a behaviour of relying on early interactions and ignoring the useful feature encountered later. It has been shown that a solution to this problem can be provided by periodically resetting the last layer of the network. According to the connections between neuroscience and RL, [12] proposes that the brain optimizes multiple cost functions to facilitate targeted learning, using efficient architectures tailored to specific cognitive tasks. It is thought that by combining both neuroscience and deep learning principles, the gap between biological and artificial intelligence can be closed according to this perspective [13]. Therefore, we can take inspiration from the relationship between reward-prediction circuits and error-processing systems to migrate plasticity loss in RL.
@@ IV-A Simulation Training and Evaluation
We evaluated the hovering flight control policy learned by the PPO agent with wind perturbations. To investigate the loss of plasticity and maintaining plasticity strategies, we trained three different RL agents, including standard PPO, L2 regularization with PPO and RECOM with L2 PPO. In Fig. 3, we have shown the average rewards for the three agents, averaged over 5 different experiments performed at 20 million timesteps.
@@
During the policy training for long periods, it was observed that the L2 and RECOM approaches produced successful results. When we analysed the change in dormant units with the L2 and RECOM approaches, we observed similar changes. However, the L2 regularization with PPO was shown to have 2.42% more dormant units than RECOM with L2 PPO. This observation led us to investigate the phenomena of dormant units in different conditions.
@@ IV-B Validation of Learned Policies
In Table III, we used the mean square error (MSE) metric to evaluate the position error of different RL agents. The standard PPO achieved a success rate of only 30% in a non-stationary environment setting, with relatively large tracking errors in three axes due to policy collapse. In contrast, L2 regularization with PPO substantially improved the success rate to 88% while also significantly reducing the overall position tracking error compared with standard PPO. Our proposed RECOM with L2 PPO method demonstrated the highest success rate at 90% and the smallest horizontal errors. There are also linear MPC approaches with sub-centimeter performance, as shown in [20]. However, the goal of this paper is not only to propose a new controller for aerial robotics applications, but also to show that a mechanism that balances rewards and losses, inspired by neuroscience and cognitive science, can evolve a policy for hovering tasks without policy collapse in long term training with non-stationary wind environment. An example of successful hovering under wind disturbance was visualized in Fig. 7.
@@ IV-C Discussions
Our experiments have carefully evaluated the proposed RECOM with L2 PPO framework, focusing on the long-term training performance of the aerial robot for the hovering task. According to the results, we show that the RECOM with L2 PPO both prevents policy collapse compared to standard PPO and organizes dormant units better than L2 regularization with PPO. While the standard PPO experienced a dramatic policy collapse after 10 million timesteps, RECOM with L2 PPO continued to attempt to maximize reward in the long-term training of the hovering task. In addition to achieve these gains, it showed that under the wind disturbance a validation performance of 90% according to the results in the RotorPy simulation environment. To better understand the contribution of rewards and losses to RL and their impact on policy stability, we considered insights from cognitive science and neuroscience.
@@
In standard RL formulations, these two signals, reward maximization and loss minimization, are usually handled indirectly (e.g., through a single scalar reward plus a training loss objective). However, by explicitly balancing the role of reward and loss in one retrospective cost term, RECOM is designed to emulate a richer feedback loop by evoking how biological systems adapt behaviour. Specifically, RECOM’s retrospective cost $\mathbb{C}_{\mathrm{ret}}$ combines the negative of cumulative reward $-\mathbb{R}[t]$ with the training loss $\mathbb{L}[t]$ into a single scalar. Therefore, it models the interplay between dopaminergic “go” signals (reward) and cortical/subcortical “caution” or “error” signals (loss). This unified feedback drives an adaptive update to the PPO learning rate, thereby aiding to preserve the agent’s plasticity over long horizons and contributing to preventing policy collapse. By periodically re-evaluating the learning rate based on recent rewards and losses, the RECOM allows the agent to both avoid policy collapse and reactivate dormant units more successfully and to adapt to changing environmental conditions throughout long-term training. Additionally, the RECOM update acts much like a negative feedback control loop, lowering the learning rate if the recent cost rises and increasing it if the cost drops. This aligns with how neural circuits modulate synaptic plasticity in response to aggregated outcomes: reward signals can boost learning rates, while substantial errors can dampen them to maintain stability.
@@ V Conclusion
In this work, we proposed a retrospective cost mechanism (RECOM) to balance rewards and losses specifically focusing on the aerial robot for hovering tasks in RotorPy with non-stationary environment scenarios. By integrating RECOM with L2 PPO, we successfully prevent policy collapse during long-term training, achieving 11.29% fewer dormant units in the policy network compared to standard PPO with L2 regularization after 20 million timesteps. The experimental results demonstrated that rewards and losses could be balanced with a retrospective mechanism inspired by the perspective of neuroscience and cognitive science.